# Is uniform expressivity too restrictive? Towards efficient expressivity of GNNs

### v1

@@ 3 Statement of the contributions
Suppose that $\xi$ is the vertex embedding computed by a GNN. We say that such GNN expresses a unary query $Q$ over a set of graphs $\mathcal{X}$ if there is a real $\epsilon<\frac{1}{2}$ such that for all graphs $G\in\mathcal{X}$ and vertices $v\in V(G)$.
$\begin{cases}\xi(G,v)\geq 1-\epsilon&\text{if $v\in Q(G)$}\\ \xi(G,v)\leq\epsilon&\text{if $v\notin Q(G)$}\end{cases}.$
@@
The following result represent the state of the art regarding the expressivity of GC2 queries by GNN with activation function $\operatorname{CReLU}$.
@@
(Barceló et al., 2020; Grohe, 2021) For every GC2 query $Q$ of size $d$, there is a GNN with activation function $\operatorname{CReLU}$, $d$ iterations and size at most $2d$ that express $Q$ over all graphs. Conversely, any (aggregation-combine) GNN expresses with activation function $\operatorname{CReLU}$ expressed a GC2 query.
@@
$\ell$ being the number of colors of the vertices in the input graphs, the family of GNNs with $\mathsf{agg}=\mathsf{sum}$, and $\mathsf{comb}(x,y)=\operatorname{ReLU}(Ax+By+C)$ (where $A\in\mathbb{N}^{\ell\times\ell}$, $B\in\mathbb{N}^{\ell\times\ell}$ and $C\in\mathbb{N}^{\ell}$) is sufficient to uniformly express all queries of GC2 uniformly. Furthermore, for each query $Q$ of depth $q$, there is a GNN of this type with at most $q$ iterations that expresses uniformly $Q$.
@@
Until now, the question about uniform expressivity has been mainly open for other activation functions such as $\tanh$ and $\operatorname{Sigmoid}$. Our contributions is two-fold. On the one hand, we show that uniform expressivity is not possible. On the other hand, we show that, despite this impossibility, a form of almost-uniform expressivity is possible.
@@ 3.1 Impossibility of Uniform Expressivity
Let $f:\mathbb{R}\rightarrow\mathbb{R}$ be a function having a finite limit $\ell\in\mathbb{R}$ in +$\infty$+ ∞. We say that $f$ converges superpolynomially to $\ell$ iff for every polynomial $P\in\mathbb{R}[X]$, $\underset{x\rightarrow+\infty}{\lim}P(x)(f(x)-\ell)\rightarrow 0$.
We say that a Pfaffian function is superpolynomial if it non constant and verifies this condition.
Our first result extends the negative result of rational GNNs by Khalife (2023) to the class of bounded superpolynomial Pfaffian GNNs.
@@
  • i)
    $Q(T,s)=1$   and   $Q(T^{\prime},s^{\prime})=0$
  • ii)
    $\lvert\xi(s^{\prime},T^{\prime})-\xi(s,T)\rvert<\epsilon$
@@
Let $\sigma\in\{\operatorname{Sigmoid},\tanh\}$. Then there is a GC2 query $Q$ such that for any GNN with activation $\sigma$, the GNN cannot express $Q$ uniformly over all graphs.
@@ 3.2 Almost-Uniform Expressivity
Our second contribution shows that for a large class of activation functions, which we call step-like activation function (see Definition 5.1 in Section 5 below) we have almost-uniform expressivity of GC2 queries.
@@
Let $\sigma$ be a step-like activation function. Then, for every GC2 query $Q$ of size $d$ and $\Delta$, there is a GNN with activation function $\sigma$, $d$ iterations and size $O(d+\log\Delta)$ that express $Q$ over all graphs with degree $\leq\Delta$. Moreover, if the parameter $\eta$ of $\sigma$ is zero, the size of the GNN can be further reduced to $O(d+\log\log\Delta)$.
@@
Let $\sigma\in\{\operatorname{Sigmoid},\tanh\}$. Then, for every GC2 query $Q$ of size $d$ and $\Delta$, there is a GNN with activation function $\sigma$, $d$ iterations and size $O(d+\log\log\Delta)$ that express $Q$ over all graphs with degree $\leq\Delta$.
We want to emphasize that the function $\log\log\Delta$ grows so slowly that for almost all theoretical and practical purposes we can consider it to be a constant function. Hence the term “almost-uniform expressivity”.
@@ 4 Superpolynomial Bounded Pfaffian GNNs: Impossibility
Let us overview why Theorem 3.2 holds, expanding all details in Appendix B. Our first result makes use of specific families of input trees that allow us to saturate the output signal of a GNN by increasing appropriately the number of vertices in certain regions of these graphs. The input trees, pictured in Figure 2 are carefully chosen in order to return distinct outputs to query of GC2. The gist of our proof consists in showing that by increasing the order of the trees, the output signals of the GNNs of the considered class can be made arbitrarily close. This is achieved by proving a stronger property on all vertices on the trees stated below. To do so, we first need the intermediate definition:
@@
where for vectors $z\in\mathbb{R}^{d}$, and functions $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $\lim_{z\rightarrow+\infty}F(z)=0$ means that for every $\epsilon>0$ there exists $\exists\beta\geq 0$, such that $\min(z_{1},\cdots,z_{d})\geq\beta$ implies $\lvert F(z)\rvert<\epsilon$.
@@
Figure 2: $T[x,m,k]$
For non-negative integers $k,m$, let $T[x,k,m]$ be the rooted tree of Figure 2. For ease of notation, if the context allows it, we will refer to $T[x,k,m]$ simply as $T$. For any non-negative integer $t$, $v\in V(T)$, let $\xi^{t}(v,T)$ be the embedding returned for node $v$ by a GNN with activation function $\sigma\in H$ after $t$ iterations. Let $M>0$ be an integer. We will prove by induction on $t$ that for any $t\in\{1,\cdots,M\}$:
@@
    $v_{t}$ is constant with respect to $x$, $k$ and $m$.
@@
    $g_{t}$ and $h_{t}$ are bounded Pfaffian functions that depend only on the combination and aggregation functions and the iteration $t$.
@@
    For every $t\in\{1,\cdots,M\}$, $\nu_{t},\eta_{t},\epsilon_{t},\epsilon^{1}_{t},\nu^{1}_{t}\in S_{3}$.
A few comments are in order. Recall that $S_{3}$ is the set of functions $\mathbb{R}^{3}\rightarrow\mathbb{R}$ such that, for any fixed first argument, the function of the second and third argument tend (collectively) superpolynomially to 00 with respect to the third argument. This asymptotic behavior will arise from our assumptions made on the Pfaffian activation functions, and will be used in our inductive argument. One immediate corollary is that we can decompose the signal at the root vertex into a main component $v_{t}$ and an error term that tends to 00 when the order of the trees to +$\infty$+ ∞. The main signal $v_{t}$ will depend only on the number of iterations, not on $x$, $m$ or $k$. In the other vertices, the main component will only be a function of $k$. This will prove for instance, that the the two outputs of the GNN on the pair $T([0,k,m],T[1,k,m])_{(k,m)\in\mathbb{N}^{2}}$ converge to the same limit, when $k$ and $m$ tend to +$\infty$+ ∞. However, fo every $k>0$ and $m>0$, $T[0,k,m]$ and $T[1,k,m]$ have distinct output at the root for the following query:
$Q(v):=\lnot(\exists^{\geq 1}y\left(E(y,v)\land(\lnot\exists^{\geq 2}vE(v,y))\right))=\forall y\left(E(y,v)\land\exists^{\geq 2}zE(z,y)\right)$
expressing that vertex $v$’s neighbors have degree at least $2$. Clearly, $Q$ is a GC2 query fo depth $4$. We can easily generalize to any GC2 query for which the root vertex of both trees have different output. In conclusion, the uniform expressivity of GNNs with superpolynomial bounded activation functions is weaker than those of the ReLUs.
@@ 5 Towards efficient non-uniform expressivity
Our second result focuses on non-uniform expressivity, where the size of the GNN is allowed to grow with the size of the input graphs. The result that we obtain holds for the wide class of activation functions defined below, and to whichever activation function can express them. This class (See Lemma C.1 in Appendix C) is characterized by a fast converge to the the linear threshold activation function
@@
Let $\sigma$ be a step-like activation function. Then, for every GC2 query $Q$ of size $d$ and $\Delta$, there is a GNN with activation function $\sigma$, $d$ iterations and size
$2d+N+\frac{2+\log(\Delta+2)}{1-\log(1-\eta)}$
parameters per iteration that express $Q$ over all graphs with degree $\leq\Delta$. Moreover, if $\eta=0$, the size of the GNN can be further reduced to
$2d+N+\log(2+\log(\Delta+2)).$
@@
We note that the fact that the size of the GNN depends on the degree $\Delta$ of the class of graphs we are working with only logarithmically guarantees that the GNN has a reasonable size for applications. Moreover, if $\eta=0$, then we have a log-log dependency, which for whichever practical setting is an almost constant function.
@@
There is a step-like activation function $\overline{\sigma}_{\arctan}$ with $\eta=0.64$, $\varepsilon=0.1$, $N=0$ and $H=1.52$ that can be expressed with a 2-layer NN with activation function $\arctan$ of size 2.
@@
There is a step-like activation function $\overline{\sigma}_{\tanh}$ with $\eta=0$, $\varepsilon=0.2$, $N=0$ and $H=2.2$ that can be expressed with a 3-layer NN with activation function $\tanh$ of size 5.
@@
There is a step-like activation function $\overline{\sigma}_{\operatorname{Sigmoid}}$ with $\eta=0$, $\varepsilon=0.2$, $N=0$ and $H=2.2$ that can be expressed with a 5-layer NN with activation function $\operatorname{Sigmoid}$ of size 9.
@@
For every GC2 query $Q$ of size $d$ and $\Delta$, there is a GNN with activation function $\tanh$, $d$ iterations and size at most $2d+1.4+0.7\log(\Delta+2)$ that express $Q$ over all graphs with degree $\leq\Delta$.
@@
For every GC2 query $Q$ of size $d$ and $\Delta$, there is a GNN with activation function $\tanh$, $d$ iterations and size at most $2d+5\log(2+\log(\Delta+2))$ that express $Q$ over all graphs with degree $\leq\Delta$.
@@
For every GC2 query $Q$ of size $d$ and $\Delta$, there is a GNN with activation function $\tanh$, $d$ iterations and size at most $2d+9\log(2+\log(\Delta+2))$ that express $Q$ over all graphs with degree $\leq\Delta$.
@@ 6.1 Queries
In this section we demonstrate how the size of the neural networks impacts the GNN’s ability to express GC2 queries via two experiments. For both of them, we consider the following queries:
$Q_{1}(v):=\lnot\left(\exists^{\geq 1}y(E(y,v)\land(\lnot\exists^{\geq 2}vE(v,y)))\right)=\forall y\left(E(y,v)\land\exists^{\geq 2}zE(z,y)\right)$
$Q_{1}(v)$ is expressing that all neigbhors of $v$ have degree at least two. Note that $Q$ is in GC2 and has depth 4444333Here, since trees are unicolored, we removed the color atomic queries for the sake of presentation. Otherwise, the depth of the query would be $5$.. In this case the input trees are unicolored.
$Q_{2}(v)=\mathsf{Red}(v)\land\left(\exists^{\geq 1}xE(x,v)\land\left(\exists^{\geq 1}vE(v,x)\land\mathsf{Blue}(v)\right)\land\left(\exists^{\geq 1}vE(v,x)\land\mathsf{Red}(v)\right)\right)$
$Q_{2}(v)$ is expressing that $v$ is red, and has a neighbor that has a red neighbor and a blue neighbor. $Q_{2}$ is in GC2 as well and has depth $7$. The vertices of the trees are colored as follows: the source and depth-one vertices are red, and only the leaves are blue.

### v2

@@ 3 Statement of the contributions
Given a unary query $Q$ and a GNN computing a vertex embedding $\xi$, we say that the GNN expresses $Q$ over a set of graphs $\mathcal{X}$ if there is a real $\epsilon<1/2$ such that for all graphs $G\in\mathcal{X}$ and vertices $v\in V(G)$,
$\begin{cases}\xi(v,G)\geq 1-\epsilon&\text{ if }Q(v,G)=1~{}\text{ and }{}\\ \xi(v,G)\leq\epsilon&\text{ if }Q(v,G)=0\end{cases}$
@@
The following result states the best known upper-bound on the number of iterations and size of GNNs with activation function $\operatorname{CReLU}$ to express GC2 uniformly.
@@
(Barceló et al., 2020; Grohe, 2021) For every GC2 query $Q$ of depth $d$, there is a GNN with activation function $\operatorname{CReLU}$, $d$ iterations and size at most $4d$ that express $Q$ over all graphs. Conversely, any GNN with activation function $\operatorname{CReLU}$ expresses a GC2 query.
@@
We can further restrict the GNN in Theorem 3.1 by requiring its combination functions to be of the form $\mathsf{comb}_{t}(x,y)=\operatorname{CReLU}(Ax+By+C)$ with $A\in\{-1,0,1\}^{d\times d}$, $B\in\{-1,0,1\}^{d\times d}$ and $C\in\mathbb{Z}^{d}$ independent of $t$.
@@
Until now, the question about uniform expressivity has been mainly open for other activation functions such as $\tanh$ and $\operatorname{Sigmoid}$. Our contributions is two-fold. On the one hand, we show that uniform expressivity is not possible. On the other hand, we show that, despite this impossibility, a form of almost-uniform expressivity is achievable.
@@ 3.1 Impossibility of Uniform Expressivity
Let $f:\mathbb{R}\rightarrow\mathbb{R}$ be a function having a finite limit $\ell\in\mathbb{R}$ in +$\infty$+ ∞. We say that $f$ converges superpolynomially to $\ell$ iff for every polynomial $P\in\mathbb{R}[X]$, $\lim_{x\to+\infty}P(x)(f(x)-\ell)=0$.
We say that a Pfaffian function is superpolynomial if it is non-constant and verifies this condition.
Our first result extends the negative result of rational GNNs to the class of bounded superpolynomial Pfaffian GNNs.
@@
  1. i)
    $Q(s,T)=1\text{ and }Q(s^{\prime},T^{\prime})=0$
  2. ii)
    $\lvert\xi(s^{\prime},T^{\prime})-\xi(s,T)\rvert<\epsilon$.
@@
Let $\sigma\in\{\operatorname{Sigmoid},\tanh\}$. There exists a GC2 query $Q$ such that no GNN with activation $\sigma$ can express $Q$ uniformly over all graphs.
@@ 3.2 Almost-Uniform Expressivity
Our second contribution shows that for a large class of activation functions, which we call step-like activation functions (see Definition 5.1 in Section 5 below) we have almost-uniform expressivity of GC2 queries.
@@
Let $\sigma$ be a step-like activation function. Then, for every GC2 query $Q$ of depth $d$ and every positive integer $\Delta$, there is a GNN with activation function $\sigma$, $d$ iterations and size $O(d\log\Delta)$ that express $Q$ over all graphs with degree $\leq\Delta$. Moreover, there exists a family of step-like activation functions such that the size of the GNN can be further reduced to $O(d\log\log\Delta)$.
@@
Let $\sigma\in\{\operatorname{Sigmoid},\tanh\}$. Then, for every GC2 query $Q$ of depth $d$ and every positive integer $\Delta$, there is a GNN with activation function $\sigma$, $d$ iterations and size $O(d\log\log\Delta)$ that express $Q$ over all graphs with degree $\leq\Delta$.
We want to emphasize that the function $\log\log\Delta$ grows so slowly that for almost all theoretical and practical purposes we can consider it to be a constant function, hence our usage of the expression “almost-uniform expressivity”.
@@ 4 Superpolynomial Bounded Pfaffian GNNs: Impossibility
Let us overview why Theorem 3.2 holds, expanding all details in Appendix B. Our first result makes use of specific families of input trees that allow us to saturate the output signal of a GNN by increasing appropriately their number of vertices in certain regions. The input trees, pictured in Figure 2 are carefully chosen in order to return distinct outputs to query of GC2. The gist of our proof consists in showing that by increasing the order of the trees in a suitable manner, the output signals of the GNNs of the considered class can be made arbitrarily close. This is achieved by proving a stronger property on all vertices on the trees stated below. To do so, we first need the intermediate definition:
@@
where for vectors $z\in\mathbb{R}^{d}$, and functions $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $\lim_{z\rightarrow+\infty}F(z)=0$ means that for every $\epsilon>0$ there exists $\beta\geq 0$, such that $\min(z_{1},\cdots,z_{d})\geq\beta$ implies $\lvert F(z)\rvert<\epsilon$.
@@
Figure 2: Tree $T[x,k,m]$ with root $s$
For non-negative integers $k$ and $m$, let $T[x,k,m]$ be the rooted tree of Figure 2. In this tree, the root $s$ has $m$ descendants $x_{i}$, having $x_{i}$ $x$ descendants for $i=1$ and $k$ descendants for $i\geq 2$. For ease of notation, whenever context allows, we will refer to $T[x,k,m]$ simply as $T$ and to any of the descendants of $x_{i}$ as $l_{i}$. For any non-negative integer $t$ and $v\in V(T)$, let $\xi^{t}(v,T)$ be the embedding returned for node $v$ by a GNN with bounded superpolynomial Pfaffian activation function $\sigma$ after $t$ iterations. Let $M>0$ be an integer. We will prove by induction on $t$ that for any $t\in\{1,\cdots,M\}$:
@@
    $v_{t}\in\mathbb{R}^{d_{t}}$ is constant with respect to $x$, $k$ and $m$.
@@
    Each coordinate of $g_{t}$ and $h_{t}$ are bounded Pfaffian functions of $k$ that depend only on the combination and aggregation functions and the iteration $t$.
@@
    Each coordinate of the functions $\nu_{t},\eta_{t},\epsilon_{t},\epsilon^{1}_{t},\nu^{1}_{t}$ are in $S_{3}$.
A few comments are in order. Recall that $S_{3}$ is the set of functions $\mathbb{R}^{3}\rightarrow\mathbb{R}$ such that, for any fixed first argument, the function of the second and third argument tend (collectively) superpolynomially to 00 with respect to the third argument. This asymptotic behavior will arise from our assumptions made on the Pfaffian activation functions, and will be used in our inductive argument. One immediate corollary is that we can decompose the signal at the root vertex into a main component $v_{t}$ and an error term that tends to 00 when the order of the trees to +$\infty$+ ∞. The main signal $v_{t}$ will depend only on the number of iterations, not on $x$, $m$ or $k$. In the other vertices, the main component will only be a function of $k$. This will prove for instance, that the the two outputs of the GNN on the pair $(T[0,k,m],T[1,k,m])_{(k,m)\in\mathbb{N}^{2}}$ converge to the same limit, when $k$ and $m$ tend to +$\infty$+ ∞. However, for every $k,m>0$, $T[0,k,m]$ and $T[1,k,m]$ have distinct output at the root for the following query:
$Q(v):=\lnot\left(\exists^{\geq 1}y\left(E(y,v)\land\left(\lnot\left(\exists^{\geq 2}v\left(E(v,y)\land\textsc{t}\right)\right)\right)\right)\right)=\forall y\left(E(y,v)\rightarrow\exists^{\geq 2}zE(z,y)\right)$
expressing that the vertex $v$’s neighbors have degree at least $2$. Note that GC2 can emulate other Boolean connectives (here $\rightarrow$) combining the ones it has, although it cannot use them directly. Clearly, $Q$ is a GC2 query of fixed depth 4. We can easily generalize to any GC2 query for which the root vertex of both trees have different output. In conclusion, the uniform expressivity of GNNs with superpolynomial bounded activation functions is weaker than those of the ReLUs.
@@ 5 Towards efficient non-uniform expressivity
Our second result focuses on non-uniform expressivity, where the size of the GNN is allowed to grow with the size of the input graphs. The result that we obtain holds for the wide class of activation functions defined below, and to any other activation function that can express them through a neural network. This class (See Lemma C.1 in Appendix C) is characterized by a fast convergence to the linear threshold activation function
@@
Recall that $\sigma^{\prime}$ denotes the derivative of $\sigma$ and $\sigma^{N}=\sigma\circ\cdots\sigma$ its $N$-th iterated composition.
@@
Remark 5.1.
Seeing $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ as a discrete dynamical system, (a) and (b) translate into 00 and $1$ being attractive fixed points; (c) says that, after $N$ iterations, every point in $\mathbb{R}\setminus(-\varepsilon,\varepsilon)$, the complement of $(-\varepsilon,\varepsilon)$, is in the attractive region of either 00 or $1$; and (d) provides a bound on a constant related to the size of these attractive regions.
@@
Let $\sigma$ be a step-like activation function. Then, for every GC2 query $Q$ of depth $d$ and every positive integer $\Delta$, there is a GNN with activation function $\sigma$, $d$ iterations and size
$\left(3+N+\frac{2+\log(\Delta+2)}{1-\log(1+\eta)}\right)d$
that express $Q$ over all graphs with degree $\leq\Delta$. Moreover, if $\eta=0$, the size of the GNN can be further reduced to
$\left(3+N+\log(2+\log(\Delta+2))\right)d.$
@@
The logarithmic dependence of the GNN’s size on the degree $\Delta$ of the considered class of graphs guarantees that the GNN has a reasonable size for applications. Moreover, if $\eta=0$, then we have a log-log dependency, which for whichever practical setting is an almost constant function.
Remark 5.3.
Our proof further shows that only the width of the combination function depends on the depth $d$ of the query, and that the number of layers remains independent of $d$.
@@
There is a step-like activation function $\overline{\sigma}_{\arctan}$ with $\eta=0.64$, $\varepsilon=0.1$, $N=0$ and $H=1.52$ that can be expressed with a 3-layer NN with activation function $\arctan$ of size 3.
@@
There is a step-like activation function $\overline{\sigma}_{\tanh}$ with $\eta=0$, $\varepsilon=0.2$, $N=0$ and $H=2.2$ that can be expressed with a 4-layer NN with activation function $\tanh$ of size 6.
@@
There is a step-like activation function $\overline{\sigma}_{\operatorname{Sigmoid}}$ with $\eta=0$, $\varepsilon=0.2$, $N=0$ and $H=2.2$ that can be expressed with a 4-layer NN with activation function $\operatorname{Sigmoid}$ of size 6.
@@
For every GC2 query $Q$ of depth $d$ and every positive integer $\Delta$, there is a GNN with activation function $\arctan$, $d$ iterations and size at most $\left(10+3.5\log(\Delta+2)\right)d$ that expresses $Q$ over all graphs with degree $\leq\Delta$.
@@
For every GC2 query $Q$ of depth $d$ and every positive integer $\Delta$, there is a GNN with activation function $\tanh$, $d$ iterations and size at most $\left(9+4\log(2+\log(\Delta+2))\right)d$ that expresses $Q$ over all graphs with degree $\leq\Delta$.
@@
For every GC2 query $Q$ of depth $d$ and every positive integer $\Delta$, there is a GNN with activation function $\operatorname{Sigmoid}$, $d$ iterations and size at most $\left(9+4\log(2+\log(\Delta+2))\right)d$ that expresses $Q$ over all graphs with degree $\leq\Delta$.
@@ 6 Numerical experiments
In this section, we demonstrate how the size of the neural networks impacts the GNN’s ability to express GC2 queries via two experiments.
@@ 6.1 Queries
In our experiments, we consider the following two queries:
$Q_{1}(v):=\lnot\left(\exists^{\geq 1}y\left(E(y,v)\land\left(\lnot\left(\exists^{\geq 2}v\left(E(v,y)\land\textsc{t}\right)\right)\right)\right)\right)=\forall y\left(E(y,v)\rightarrow\exists^{\geq 2}zE(z,y)\right)$
and
$Q_{2}(v)=\mathsf{Red}(v)\land\left(\exists^{\geq 1}x\left(E(x,v)\land\left(\exists^{\geq 1}v\left(E(v,x)\land\mathsf{Blue}(v)\right)\right)\land\left(\exists^{\geq 1}v\left(E(v,x)\land\mathsf{Red}(v)\right)\right)\right)\right).$
One can see that $Q_{1}(v)$ expresses that all neighbors of $v$ have degree at least two, and $Q_{2}(v)$ that $v$ is red and it has a neighbor that has a red neighbor and a blue neighbor. Moreover, $Q_{1}$ has depth $4$ and $Q_{2}$ has depth $7$. Note that $Q_{1}$ applies even to graphs with any number of colors, while $Q_{2}$ applies to graph having at least two colors.